OpenStudy (anonymous):

soybean meal is 12% protein; cornmeal is 6% protein. How many lbs of each should be mixed together in order to get a 240 lb mixture that is 7% protein? How many lbs of cornmeal? how many lbs of soybean?

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

waittttt.

OpenStudy (saifoo.khan):

those are inn answers.

OpenStudy (saifoo.khan):

inccrt.

OpenStudy (anonymous):

ok.

OpenStudy (saifoo.khan):

Let s be the number of pounds of soybean meal and c be the number of pounds of cornmeal. Make a system of equations. s + c = 240 0.12s + 0.06c = 0.07(240) or 16.8 Solve by substitution. s + c = 240 s = 240 - c 0.12s + 0.06c = 16.8 (substitute 240 - c for s) 0.12(240 - c) + 0.06c = 16.8 28.8 - 0.12c + 0.06c = 16.8 28.8 - 0.06c = 16.8 -0.06c = -12 c = 200 s = 240 - c = 240 - 200 = 40 ANSWER: 200 pounds of cornmeal and 40 pounds of soybean meal

OpenStudy (saifoo.khan):

here u go!

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

You Welcome!

OpenStudy (anonymous):

you have been a life saver!

OpenStudy (saifoo.khan):

hehe, Thnks! :)

OpenStudy (anonymous):

I am about to submit my assignment...I will let you know how i did in a couple minutes